# Towards Generalizable Autonomous Penetration Testing via Domain Randomization and Meta-Reinforcement Learning

### v1

@@ VI Discussion
VII Conclusion
In this paper, we present GAP, a framework for training generalizable pentesting agents capable of drawing inferences from one to another. To achieve this, we introduce a Real-to-Sim-to-Real pipeline within GAP, which integrates two critical methods: LLM-powered domain randomization and Meta-RL training. The preliminary evaluations in this paper demonstrate that GAP allows pentesting agents for end-to-end policy learning in unknown environments, bridging the generalization gap for zero-shot policy transfer in similar environments, and facilitating rapid policy adaption in dissimilar environments. GAP fills a critical research gap as the first framework of its kind. By introducing GAP, we not only contribute to the field of autonomous pentesting but also provide a feasible framework that can be used in other real-world applications.

### v2

@@ 7 Discussion
8 Conclusion
In this paper, we present GAP, an autonomous pentesting framework for efficient policy training in realistic environments and for training generalizable agents capable of drawing inferences about other cases from one instance – a key to the broad application of autonomous pentesting agents. To achieve this, GAP introduces a Real-to-Sim-to-Real pipeline that (a) enables end-to-end policy learning in unknown realistic environments while constructing realistic simulation analogs; (b) improves agents’ generalization ability by leveraging domain randomization and meta-RL learning. The preliminary evaluations demonstrate that GAP allows pentesting agents for end-to-end policy learning in realistic environments, bridging the generalization gap for zero-shot policy transfer in similar environments, and facilitating rapid policy adaptation in dissimilar environments. GAP fills a critical research gap as the first framework of its kind. By introducing GAP, we not only contribute to the field of autonomous pentesting but also provide a feasible framework that may be used in other applications.